Find the percentage of calcium in calcium carbonate (CaCO3)

Grade:10

2 Answers

The Hounds Of Justice
5 years ago
Molecular mass of CaCO3 = 40+12+(16*3) = 100u
Atomic mass of Ca= 40
Percentage of Ca= (40/100)*100= 40%
Arun
5 years ago
Write out the formula first: CaCO3. Next, find the mass of each of them by using the PToE (below the element symbol). 

Ca - 40.08g/mol 
C - 12.01g/mol 
O - 3*16.00g/mol = 48.00g/mol 

Add them all together to find the molar mass: 
40.08 
12.01 
+ 48.00 
--------------- 
100.09g/mol 

So you have the molar mass. Now you want to find the percentage of calcium, so get the mass from above and divide it by the whole molar mass. 

40.08g/mol divided by 100.09g/mol (times this by 100 to get a percent). 

Your answer would be 40.04%. Hope this helps!